1. At a distance of R = 8500 km from the center of the Earth an artificial satellite moves at a speed V = 6.8 km / s. Calculate the force of gravitational pull exerted by the Earth on a satellite of mass m = 10t


1
Expert's answer
2021-01-19T11:58:53-0500

"F=mv^2/r=10000\cdot6800^2/8500000=54400(N)"